# A new understanding of Einstein-Rosen bridges

### v1

@@ Abstract
The formulation of quantum field theory in Minkowski spacetime, which emerges from the unification of special relativity and quantum mechanics, is based on treating time as a parameter, assuming a fixed arrow of time, and requiring that field operators commute for spacelike distances. This procedure is questioned here in the context of quantum field theory in curved spacetime (QFTCS). In 1935, Einstein and Rosen (ER), in their seminal paper [1] proposed that ”a particle in the physical Universe has to be described by mathematical bridges connecting two sheets of spacetime” which involved two arrows of time. Recently proposed direct-sum quantum theory reconciles this ER’s vision by introducing geometric superselection sectors associated with the regions of spacetime related by discrete transformations. We further establish that the quantum effects at gravitational horizons involve the physics of quantum inverted harmonic oscillators that have phase space horizons. This new understanding of the ER bridges is not related to classical wormholes, it addresses the original ER puzzle and promises a unitary description of QFTCS, along with observer complementarity. Furthermore, we present compelling evidence for our new understanding of ER bridges in the form of large-scale parity asymmetric features in the cosmic microwave background, which is statistically 650 times stronger than the standard scale-invariant power spectrum from the typical understanding of inflationary quantum fluctuations when compared with the posterior probabilities associated with the model given the data. We finally discuss the implications of this new understanding in combining gravity and quantum mechanics.
@@ 1 Introduction
We discuss the relations between ER bridges and analogous proposals in different contexts with the recently developed framework of direct-sum quantum theory and its applications to early Universe cosmology, and BH physics [23, 24, 25, 26, 27, 28, 29]. Direct-sum quantum theory is based on the discrete spacetime (such as parity ($\mathcal{P}$) and time reversal ($\mathcal{T}$)) (a)symmetries of the physical system to formulate a description of the quantum state as a direct-sum of components corresponding to (geometric) superselection sectors (SSS) of Hilbert space. Geometric SSS are Hilbert spaces that describe quantum states corresponding to regions of physical space related by discrete spacetime transformations. If a Hilbert space is a direct-sum of geometric SSS, a state vector in that Hilbert space becomes a direct-sum of components, each corresponding to geometric SSS. The same applies to operators in the Hilbert space. This is called the geometric superselection rule. We show that the ”direct-sum” is the mathematical bridge that matches the expectations of ER bridges in describing one physical world with two arrows of time. The two arrows of time here operate at the parity conjugate regions of physical space embedded with the geometric construction of SSS. This framework restores unitarity in curved spacetime, and it is tested against the latest observations of the cosmic microwave background from the Planck satellite data.
@@ 2 Quantum fields in Minkowski spacetime and examples of (inverted) harmonic oscillator physics
  • Define a positive energy state with respect to the arrow of time: $|\Psi\rangle_{t}=e^{-i\mathcal{E}t_{p}}|\Psi\rangle_{0}$
@@
which is again an infinite collection of harmonic oscillators in 3 spatial dimensions. What if we consider the case $\omega\to i\omega?$ (i.e., $\tilde{\omega}^{2}=-\omega^{2}>0$ in (9) and correspondingly $m\to im?$ (i.e., $\mu^{2}=-m^{2}>0$ in (7). This would turn the harmonic oscillator into an IHO and the corresponding field to tachyonic field (i.e., infinite collection of coupled IHOs)
@@ 2.1 Time translations and Killing vector of Minkowski spacetime:
where $J^{0}$ is the zeroth component of conserved current $J^{\mu}$ and $T^{00}$ is the $00$th component of energy-momentum tensor $T^{\mu\nu}$. The crucial point we must notice here is that the Killing vector (17) retain its property (16) under the time reversal transformation $t_{p}\to-t_{p}$. Thus, the conserved energy $\mathbb{E}$ remain the same under $t_{p}\to-t_{p}$. This is not surprising as the time reversal operation is (discrete) symmetry (2) of the Minkowski manifold (1).
@@ 3 Review of inverted harmonic oscillator (IHO), quantization and Riemann Hypothesis
As we have seen in the previous section, Higgs Mexican hat potential (13) indicates the first appearance of IHO in SM physics. As we will further demonstrate in the later sections, in the context of early Universe cosmology, the inflationary quantum fluctuations, which manifestly appear in terms of canonical quantum variables (known as Mukhanov-Sasaki variables) [36], can be understood through QFT in terms of inverted harmonic oscillators (IHOs). Even in the context of Black Hole physics, IHOs are found to be the fundamental building blocks to describe Hawking radiation [37, 38, 35]. The role of IHOs even extends to the Rindler spacetime and also in the context of the quantum Hall effect, molecular physics, and even in biophysics (See [35, 39] and references therein). Moreover, the quantum aspects of the IHO manifest in relation to the nontrivial zeros of the Riemann zeta function [40]. This section provides a comprehensive brief review of IHO physics, quantization, and its non-trivial relation to the Riemann zeta function. In particular, we emphasize how foundational questions regarding the quantization of the IHO are intimately connected to the challenge of establishing its energy spectrum as corresponding to the nontrivial zeros of the Riemann zeta function. The elements discussed in this section, especially the wave functions and phase space regions of the IHO, are closely tied to the physics of quantum fields in curved spacetime, which we will develop in the sections that follow.
@@
which is known as the Berry-Keating Hamiltonian [22] whose motion equations give the following solutions
@@
As a consequence of Heisenberg’s uncertainty principle, we have
@@
The position $Q$ and momentum $P$ wavefunctions of the IHO Hamiltonian operator (for the region $Q>0$ and $E<0$) are [41, 22, 38]
@@
  • The wavefunction $\Psi\left(Q\right)$ is also an Eigen function of Weyl reflected Laplace-Beltrami operator
    $L_{R}=-Q^{2}\partial_{Q}^{2}-2Q\partial_{Q}=\left(\frac{1}{2}-\frac{i\hat{H}_{iho}}{\hbar\omega}\right)\left(\frac{1}{2}+\frac{i\hat{H}_{iho}}{\hbar\omega}\right)$ (32)
    with positive definite Eigenvalues $\left(\frac{1}{4}+\frac{E^{2}}{\hbar^{2}}\right)$.
  • BK proposes identifying the discrete set of points in phase space, which are $\left(Q,\,P\right)$, and $\left(-Q,\,-P\right)$. This is very much similar to the antipodal identification in dS spacetime proposed by Schrödinger and the one of ’t Hooft in the context of Schwarzschild spacetime [3, 21]. As we will discuss in the later sections, the antipodal identification is similar to the ER’s mathematical bridge. Thus, what BK proposes is another ”mathematical bridge” to join the IHO’s phase space regions with opposite arrows of time.
@@ 4.1 QFTCS in BH spacetime
$T=\frac{U+V}{2},\,X=\frac{V-U}{2}$ (45)
@@ 4.2 QFTCS in de Sitter spacetime
where $R_{dS}$ is the curvature scalar of de Sitter space and $\tau=\int\frac{dt}{a}$ being the conformal time. One thumb rule in physics is to make use of symmetries. The origin of physical conundrums often occurs, throwing away any symmetries by hand. Similar to SBH, dS spacetime, too (58), describes the physical world with two possible arrows of time given by
@@
In this picture, the Universe evolves from contraction to expansion (bounce), which is compatible with two arrows of time
${\rm Contraction\,to\,Expansion}\implies\begin{cases}t:-\infty\to+\infty,\quad H>0,\quad\left(\tau:\frac{\pi}{2|H|}\to-\frac{\pi}{2|H|}\right)\\ t:+\infty\to-\infty,\quad H<0,\quad\left(\tau:-\frac{\pi}{2|H|}\to\frac{\pi}{2|H|}\right)\end{cases}$ (66)
@@ 4.2.1 Occurrence of IHO physics in de Sitter spacetime:
$\mu_{eff}^{2}=H^{2}\cos(2H\tau)>0$ (71)
Extending this to a massive KG field we will get
$\mu_{eff}^{2}$ $=H^{2}\cos(2H\tau)-m^{2}a^{2}$ (72)
$=\frac{2H^{2}}{a^{2}}-H^{2}-m^{2}a^{2}$
@@ 4.3.1 IHO analogy with Rindler spacetime
$U_{R}=\pm\frac{1}{a}e^{\xi-t_{R}},\quad V_{R}=\pm\frac{1}{a}e^{\xi+t_{R}}$ (81)
@@ 4.5 Unitarity of standard QFTCS: echoing ideas from Einstein-Rosen, Schrödinger and ’t Hooft
In all the spacetime manifolds we discussed in the previous sections, which are BH, Rindler, de Sitter, and inflationary spacetimes, we witnessed identical degenerate regions of spacetimes just related by discrete transformations. Historically, QFT in curved spacetime [72] has always been carried out by choosing one of the regions as physical to the real world description. For example, in the context of BH spacetime, Hawking’s assumption [4] is only the regions U<0,V>0U<0,\,V>0 and U>0,V<0U>0,\,V<0 are considered to be physical. In the context of dS and inflationary spacetimes, the common practice is to choose $\tau<0$ for an expanding Universe [68]. This procedure not only breaks the symmetries of the manifolds but also the unitarity and observer complementarity.
@@
This leads to a local observer accessing only part of the entanglement (mixed state) and it is violation of unitarity [14, 67]. The emergence of (89) is deeply connected to the treatment of spacetime regions bounded by gravitational horizons as open quantum systems. This is known as the central dogma in cosmology and black hole physics [47, 61]. Furthermore, unitarity in standard QFTCS is broken also because of time reversal symmetries in dS and BH spacetimes allow us to describe physical world with two arrows of time for both inside and outside the horizon. As we discussed earlier, this has led to the conundrum of entangled Universes separated by spacelike distances. All the investigations over the decades have admitted the inevitability of unitarity loss in curved spacetime unless a new physics is built from an unknown theory of quantum gravity at the Planck scales. Even then, the particle description problem initiated by ER remains unsolved. Thus, QFTCS remains one of the deepest problems in theoretical physics that impedes the progress in achieving quantum gravity [10].
@@ 5.1 Direct-sum quantum mechanics
We note that $\mathcal{P}\mathcal{T}$ operations remain the same irrespective of any coordinate translations. Thus, one can shift the origin $x=0$, but the direct-sum quantum theory is unaltered. Once we divide the quantum state by the above direct-sum operation into sectorial Hilbert space, we can still perform individually time reversal and parity operations in $\mathcal{H}_{\pm}$. The $\mathcal{T}$ operation in each SSS turns the positive energy states to negative ones and changes the direction of momenta, whereas the $\mathcal{P}$ operation changes only the direction of momenta.
@@ 5.3 Direct-sum quantization of Berry and Keating IHO
which describes the evolution of positive and negative energy quantum states. Due to the doubly degenerate regions II (IIII) and IIII (IVIV) related by (26) in the separatrix phase space (Fig. 1), the states $|\Psi_{iho}^{\pm E}\rangle$ would then be governed by
@@
where the Hamiltonians correspond to each region are functions of corresponding position and momentum operators (126). Since the positive and negative energy regions are related by the transformations (27), the position and momentum wave functions in the region II (IIII) and IIIIII (IVIV) swap with each other (i.e., Fourier transform in the region II (IIII) becomes inverse Fourier transform in the region IIIIII (IVIV) and vice versa). Working out (128) we obtain
@@ 6 Quantum ER bridges and unitarity in Schwarschild, de Sitter, and Rindler spacetimes
In this section, we delve deeper into the consequences of the direct-sum quantum theory for quantum field descriptions in curved spacetime, with particular attention to Rindler spacetime, which represents Minkowski spacetime from the perspective of a uniformly accelerated observer[25]. Remarkably, Einstein and Rosen’s work also touches upon quantum physics in Rindler like spacetimes, emphasizing the necessity of constructing ”mathematical bridges” to unify seemingly disconnected regions [1].

### v2

@@ Abstract
The formulation of quantum field theory in Minkowski spacetime, which emerges from the unification of special relativity and quantum mechanics, is based on treating time as a parameter, assuming a fixed arrow of time, and requiring that field operators commute for spacelike distances. This procedure is questioned here in the context of quantum field theory in curved spacetime (QFTCS). In 1935, Einstein and Rosen (ER), in their seminal paper (Einstein and Rosen 1935 Phys. Rev. 48 73–77) proposed that ”a particle in the physical Universe has to be described by mathematical bridges connecting two sheets of spacetime” which involved two arrows of time. Recently proposed direct-sum quantum theory reconciles this ER’s vision by introducing geometric superselection sectors associated with the regions of spacetime related by discrete transformations. We further establish that the quantum effects at gravitational horizons involve the physics of quantum inverted harmonic oscillators that have phase space horizons. This new understanding of the ER bridges is not related to classical wormholes, it addresses the original ER puzzle and promises a unitary description of QFTCS, along with observer complementarity. Furthermore, we present compelling evidence for our new understanding of ER bridges in the form of large-scale parity asymmetric features in the cosmic microwave background, which is statistically 650 times stronger than the standard scale-invariant power spectrum from the typical understanding of inflationary quantum fluctuations when compared with the posterior probabilities associated with the model given the data. We finally discuss the implications of this new understanding in combining gravity and quantum mechanics.
@@ 1 Introduction
We discuss the relations between ER bridges and analogous proposals in different contexts with the recently developed framework of direct-sum quantum theory and its applications to early Universe cosmology, and BH physics [23, 24, 25, 26, 27, 28, 29]. Direct-sum quantum theory is based on the discrete spacetime (such as parity ($\mathcal{P}$) and time reversal ($\mathcal{T}$)) (a)symmetries of the physical system to formulate a description of the quantum state as a direct-sum of components corresponding to (geometric) superselection sectors (SSS) of Hilbert space. Geometric SSS are Hilbert spaces that describe quantum states corresponding to regions of physical space related by discrete spacetime transformations. If a Hilbert space is a direct-sum of geometric SSS, a state vector in that Hilbert space becomes a direct-sum of components, corresponding to geometric SSS. The same applies to operators in the Hilbert space. This is called the geometric superselection rule. We show that the ”direct-sum” is the mathematical bridge that matches the expectations of ER bridges in describing one physical world with two arrows of time. The two arrows of time here operate at the parity conjugate regions of physical space embedded with the geometric construction of SSS. This framework restores unitarity in curved spacetime, and it is tested against the latest observations of the cosmic microwave background from the Planck satellite data.
@@ 2 Quantum fields in Minkowski spacetime and examples of (inverted) harmonic oscillator physics
  • Define a positive energy ($\mathcal{E}>0$) state with respect to the arrow of time: $|\Psi\rangle_{t}=e^{-i\mathcal{E}t_{p}}|\Psi\rangle_{0}$
@@
which is again an infinite collection of harmonic oscillators in 3 spatial dimensions. What if we consider the case $\omega\to i\omega?$ i.e., $\tilde{\omega}^{2}=-\omega^{2}>0$ in (9) and correspondingly $m\to im?$ (i.e., $\mu^{2}=-m^{2}>0$ in (7). This would turn the harmonic oscillator into an IHO and the corresponding field into a tachyonic field (i.e., an infinite collection of coupled IHOs)
@@ 2.1 Time translations and Killing vector of Minkowski spacetime:
where $J^{0}$ is the zeroth component of conserved current $J^{\mu}$ and $T^{00}$ is the $00$th component of energy-momentum tensor $T^{\mu\nu}$. The crucial point we must notice here is that the Killing vector (17) retains its property (16) under the time reversal transformation $t_{p}\to-t_{p}$. Thus, the conserved energy $\mathbb{E}$ remains the same under $t_{p}\to-t_{p}$. This is not surprising as the time reversal operation is (discrete) symmetry (2) of the Minkowski manifold (1).
@@ 3 Review of inverted harmonic oscillator (IHO), quantization and Riemann Hypothesis
As we have seen in the previous section, the Higgs Mexican hat potential (13) indicates the first appearance of IHO in SM physics. As we will further demonstrate in the later sections, in the context of early Universe cosmology, the inflationary quantum fluctuations, which manifestly appear in terms of canonical quantum variables (known as Mukhanov-Sasaki variables) [36], can be understood through QFT in terms of inverted harmonic oscillators (IHOs). Even in the context of Black Hole physics, IHOs are found to be the fundamental building blocks to describe Hawking radiation [37, 38, 35]. The role of IHOs even extends to the Rindler spacetime and also in the context of the quantum Hall effect, molecular physics, and even in biophysics (See [35, 39] and references therein). Moreover, the quantum aspects of the IHO manifest in relation to the nontrivial zeros of the Riemann zeta function [40]. This section provides a comprehensive brief review of IHO physics, quantization, and its non-trivial relation to the Riemann zeta function. In particular, we emphasize how foundational questions regarding the quantization of the IHO are intimately connected to the challenge of establishing its energy spectrum as corresponding to the nontrivial zeros of the Riemann zeta function. The elements discussed in this section, especially the wave functions and phase space regions of the IHO, are closely tied to the physics of quantum fields in curved spacetime, which we will develop in the sections that follow.
@@
which is known as the Berry-Keating Hamiltonian [22] whose equations of motion give the following solutions
@@
As a consequence of Heisenberg’s uncertainty principle, we have (restoring here the factors of $\hbar$)
@@
The position $Q$ and momentum $P$ wavefunctions of the IHO Hamiltonian operator (for the region $Q>0$ and $E<0$) are (in the units of setting $\omega=1$) [41, 22, 38]
@@
  • The wavefunction $\Psi\left(Q\right)$ is also an Eigen function of the Weyl reflected Laplace-Beltrami operator (expressed here in the units of setting $\omega=1$)
    $L_{R}=-Q^{2}\partial_{Q}^{2}-2Q\partial_{Q}=\left(\frac{1}{2}-\frac{i\hat{H}_{iho}}{\hbar}\right)\left(\frac{1}{2}+\frac{i\hat{H}_{iho}}{\hbar}\right)$ (32)
    with positive definite Eigenvalues $\left(\frac{1}{4}+\frac{E^{2}}{4\hbar^{2}}\right)$.
  • BK proposes identifying the discrete set of points in phase space, which are $\left(Q,\,P\right)$, and $\left(-Q,\,-P\right)$. This is very much similar to the antipodal identification in de Sitter spacetime proposed by Schrödinger and the one of ’t Hooft in the context of Schwarzschild spacetime [3, 21]. As we will discuss in the later sections, the antipodal identification is similar to the ER’s mathematical bridge. Thus, what BK proposes is another ”mathematical bridge” to join the IHO’s phase space regions with opposite arrows of time.
@@ 4.1 QFTCS in BH spacetime
$T=\frac{U+V}{2\sqrt{e}},\,X=\frac{V-U}{2\sqrt{e}}$ (45)
@@ 4.2 QFTCS in de Sitter spacetime
where $R_{dS}$ is the curvature scalar of de Sitter space and $\tau=\int\frac{dt}{a}$ being the conformal time and $H=\frac{\dot{a}}{a}$ is the Hubble parameter. One thumb rule in physics is to make use of symmetries. The origin of physical conundrums often occurs, throwing away any symmetries by hand. Similar to SBH, dS spacetime, too (58), describes the physical world with two possible arrows of time given by
@@
In this picture, the Universe evolves from contraction to expansion (bounce), which is compatible with two arrows of time. Such a bounce can arise from a closed FLRW cloud evolving from contraction to expansion through a degenerate quantum ground state of constant energy density (see [57]).
${\rm Contraction\,to\,Expansion}\implies\begin{cases}t:-\infty\to+\infty,\quad H>0,\quad\left(\tau:-\frac{\pi}{2|H|}\to\frac{\pi}{2|H|}\right)\\ t:+\infty\to-\infty,\quad H<0,\quad\left(\tau:\frac{\pi}{2|H|}\to-\frac{\pi}{2|H|}\right)\end{cases}$ (66)
It is worth noting that the two arrows of time (66) of a bouncing Universe [57] fit nicely with James Hartle’s proposal [58] of the Universe emerging from a time-symmetric quantum phase, which is a concept well-rooted in the no boundary proposal in quantum cosmology [59, 60].
@@ 4.2.1 Occurrence of IHO physics in de Sitter spacetime:
$\mu_{eff}^{2}=\frac{a^{\prime\prime}}{a}=H^{2}\!\left(2\sec^{2}(H\tau)-1\right)=H^{2}\!\left(1+2\tan^{2}(H\tau)\right)>0$ (71)
Extending this to a massive KG field, we will get
$\mu_{\mathrm{eff}}^{2}(\tau)=\frac{a^{\prime\prime}}{a}-m^{2}a^{2}=H^{2}\!\left(1+2\tan^{2}(H\tau)\right)-m^{2}a^{2}$ (72)
@@ 4.3.1 IHO analogy with Rindler spacetime
$U_{R}=\pm\frac{1}{a}e^{a\xi-at_{R}},\quad V_{R}=\pm\frac{1}{a}e^{a\xi+at_{R}}$ (81)
@@ 4.5 Unitarity of standard QFTCS: echoing ideas from Einstein-Rosen, Schrödinger and ’t Hooft
In all the spacetime manifolds we discussed in the previous sections, which are BH, Rindler, de Sitter, and inflationary spacetimes, we witnessed identical degenerate regions of spacetimes just related by discrete transformations. Historically, QFT in curved spacetime [76] has always been carried out by choosing one of the regions as physical to the real world description. For example, in the context of BH spacetime, Hawking’s assumption [4] is only the regions U<0,V>0U<0,\,V>0 and U>0,V>0U>0,\,V>0 are considered to be physical. In the context of dS and inflationary spacetimes, the common practice is to choose $\tau<0$ for an expanding Universe [72]. This procedure not only breaks the symmetries of the manifolds but also the unitarity and observer complementarity.
@@
This leads to a local observer accessing only part of the entanglement (mixed state) and it is violation of unitarity [14, 71]. The emergence of (89) is deeply connected to the treatment of spacetime regions bounded by gravitational horizons as open quantum systems. This is known as the central dogma in cosmology and black hole physics [47, 65]. Furthermore, unitarity in standard QFTCS is broken also because of time reversal symmetries in dS and BH spacetimes allow us to describe physical world with two arrows of time for both inside and outside the horizon. As we discussed earlier, this has led to the conundrum of entangled Universes separated by spacelike distances. All the investigations over the decades have admitted the inevitability of unitarity loss in curved spacetime unless a new physics is built from an unknown theory of quantum gravity at the Planck scales. Even then, the particle description problem initiated by ER remained unsolved. Thus, QFTCS remains one of the deepest problems in theoretical physics that impedes the progress in achieving quantum gravity [10].
@@ 5.1 Direct-sum quantum mechanics
We note that $\mathcal{P}\mathcal{T}$ operations remain the same irrespective of any coordinate translations. Thus, one can shift the origin $x=0$, but the direct-sum quantum theory is unaltered. Once we divide the quantum state by the above direct-sum operation into sectoral Hilbert space, we can still perform time reversal and parity operations individually in $\mathcal{H}_{\pm}$. The $\mathcal{T}$ operation in each SSS turns the positive energy states to negative ones and changes the direction of momenta, whereas the $\mathcal{P}$ operation changes only the direction of momenta. In Fig. 3 we illustrate how the quantum harmonic oscillator can be viewed in a new way with our direct-sum quantum theory.
@@ 5.3 Direct-sum quantization of Berry and Keating IHO
which describes the evolution of positive and negative energy quantum states. Due to the doubly degenerate regions II (IIIIII) and IIII (IVIV) related by (26) in the separatrix phase space (Fig. 1), the states $|\Psi_{iho}^{\pm E}\rangle$ would then be governed by
@@
where the Hamiltonians correspond to each region are functions of corresponding position and momentum operators (126). Since the positive and negative energy regions are related by the transformations (27), the position and momentum wave functions in the region II (IIII) and IIIIII (IVIV) swap with each other (i.e., Fourier transform in the region II (IIII) becomes inverse Fourier transform in the region IIIIII (IVIV) and vice versa). Working out (128) we obtain (in the units of setting $\omega=1$)
@@ 6 Quantum ER bridges and unitarity in Schwarschild, de Sitter, and Rindler spacetimes
In this section, we delve deeper into the consequences of the direct-sum quantum theory for quantum field descriptions in curved spacetime, with particular attention to Rindler spacetime, which represents Minkowski spacetime from the perspective of a uniformly accelerated observer [25]. The foundations of DQFT are extended to Rindler spacetime, and a renewed understanding of the Reeh-Schlieder theorem and entanglement structure in flat spacetime is derived in detail in [25]. Remarkably, Einstein and Rosen’s work also touches upon quantum physics in Rindler-like spacetimes, emphasizing the necessity of constructing ”mathematical bridges” to unify seemingly disconnected regions [1].